OpenStudy (quickstudent):

Is this correct? question: Becky is growing a garden in her backyard. It is in the shape of a rectangle 6 feet wide by 3 feet long. Her husband said that he can grow the same amount of crop in a third of the space. What are the dimensions of his garden? my answer: wide = 1/3 * 6 = 2 long = 1/3 * 3 = 1 So, the dimensions are 2 feet by 1 feet.

OpenStudy (phi):

He uses 1/3 the area is 2 sq ft 1/3 the area of 18 sq ft ?

OpenStudy (jiteshmeghwal9):

let the dimensions be x & y xy=1/3(6*3) 6/x=3/y

OpenStudy (jiteshmeghwal9):

find x & y

OpenStudy (phi):

to answer this question, I would first figure out the area of Becky's garden 3*6 = 18 then I would divide that by 3 to get 18/3 = 6 sq. ft then I would look for an answer choice that has 6 sq ft as the answer

OpenStudy (quickstudent):

alright, here's my new answer: 6 * 3 = 18 1/3 * 18 = 6 If the area is 6 square feet, then the dimensions would be 3 * 2.

OpenStudy (phi):

They do not say the smaller plot is "similar" to the larger plot. But if we wanted the dimensions of a similar plot, we would divide the L and W by sqr(3) in this case we would get 6/sqr(3) * 3/sqr(3) which can be written as 2 sqr(3) * sqr(3) if it does not have to be similar, there are lots (infinite) number of choices. 2*3 works so does 1*6 and so on.

OpenStudy (quickstudent):

Yeah I know, but it did ask what were the dimensions, so I used that.

OpenStudy (phi):

They may want the smaller plot to be similar, so be sure you can give the dimensions for that case. The idea is if the *area* is smaller by 1/3 , then each "length" is 1/sqr(3) smaller